OpenStudy (anonymous):

For the integral of f(x)= ((t^2-4)/(1+t^4)dt) from 0 to x how should you classify the relative maximum and minimum of f(x)

OpenStudy (mimi_x3):

\[\large \int\limits\frac{t^{2}-4}{1+t^{4}} dt\] Trig sub might work.. \(x^2=tan\theta\)

OpenStudy (mimi_x3):

wait \(t^2=tan\theta\) typo

OpenStudy (anonymous):

@Beckylynn u mean to find x that makes f(x) maximum or minimum?

OpenStudy (mimi_x3):

woops then you differentiate? lol

OpenStudy (anonymous):

why Beckylynn used word classify ?

OpenStudy (mimi_x3):

then what do you do ? Beckylynn please respond we need clarifications.